Question 9

Last year, you bought a bond with face value $1000, maturity 15 years, coupon rate of 5.5% per year payable semi-annually and yield to maturity of 7% per year. Currently the bond sells for $900. How much would be your total yield if you sell this bond today?

13.71%

10.78%.

(17.84%)

(15.79%)

please don't copy answers from other posts... Double check the answer.. I need it now.. thank you
Answer #1

Coupon Rate =5.5%
coupon =5.5%*1000/2 = 27.50
Number of Periods = 15*2 = 30
Semi annual YTM =7%/2 =3.5%
Price of bond Last Year =PV of coupons + Pv of par Value =27.50*(1-(1+3.5%)^30+1000/(1+3.5%)^30 =862.06

Price of Bond This Year = 900

Total Yield of Bond today =( Price Today - Price Last Year + 2* Coupons)/Price Last Year =(900 -862.06 +2* 27.50)/862.06
=10.78%
(Option b is correct option)